TT2 Meeting (Switching magnets and transport)

→ Europe/Zurich
865/2-C19 (CERN)

865/2-C19

CERN

10
Show room on map

Present: J. Coupard, S. Pittet, K. Hanke, H. Damerau, G. Sterbini, L. de Mallac, C. Bertone

K. Hanke and H. Damerau informed that the re-baselining was sent out on last Friday. The TT2 activities are not yet included and will be included in middle of this year. 

D. Bodart could not join the meeting. He sent an email summarising the relevant aspects of the TT2 switching magnets.

"[...] The two magnets to be installed are part of our 8 available magnet spares, that will be certified starting in February. However, the magnets shall be made ready for LS2 as far as understood.

As explained by C. Bertone, transport costs are estimated to be 10 kCHF, vacuum work consisting of removal and re-installation of the chamber and leak checks are estimated at 2 days or about 2 kCHF (no cutting of the chambers), survey is 1 day maximum or about 1 kCHF, magnet team is about ½ day or 500 CHF. The magnet certifications is usually covered by the operation budgets.

Globally the 20 kCHF initially requested covers largely the intervention costs.

This is a typical activity that we are used to perform during machine accesses, and which is usually covered in the coordination meeting of the machines with Simon."

J. Coupard asked if the TT2 magnet activities in LIU-PS include only the two switching dipoles. K. Hanke answered positively.

After discussion it was decided to split the requested budget (20 kCHF) in the transport WP held by C. Bertone (10 kCHF)  and in magnet WP held by D. Bodart (10 kCHF). The magnet WP will pay the survey and vacuum activities related to the installation.

C. Bertone informed that the total budget requested in the new baseline for the TT2 WP is 30 kCHF (20 kCHF for the the TT2 EPC installation and 10 kCHF for the TT2 magnet installation).

H. Damerau asked if the gateways for the power converters are included in the EPC WP. S. Pittet answer positively and explained that EPC will buy them from CO.  Concerning the possibility of implementing of the MTG flag in the converters,  S. Pittet informed that there is not a clear answer yet but a proposal will follow soon. 

K. Hanke informed that discussions are ongoing to move in LIU-PS also the consolidation of the nToF EPC converters. The preference solution of the WP holders is to import nToF in LIU-PS.

Next meeting will be held on Monday 8 Feb at 9h00 to summarise the TT2 Task Force activities.

 

 

 

There are minutes attached to this event. Show them.